OpenStudy (anonymous):

Solve the linear equation:

OpenStudy (anonymous):

\[-6x+6-7(x+1)=6x-7\]

OpenStudy (anonymous):

First distribute the 7 into the brackets. let me know what you get.

OpenStudy (anonymous):

-6x+6-7x-7=6x-7?

OpenStudy (anonymous):

correct. now, we need to get all the terms containing x on the same side of the equality. I suggest adding 6x to each side. what do you get?

OpenStudy (anonymous):

Or sorry... subtract 6x on each side.

OpenStudy (anonymous):

less work that way

OpenStudy (anonymous):

-6x-7x-7=-12x-7

OpenStudy (anonymous):

not quite. we should get:\[-6x+6-7x-7-6x=6x-6x-7\]Notice that 6x-6x=0 on the right, giving:\[-6x+6-7x-7-6x=7\]

OpenStudy (anonymous):

damn typos today: -6x+6-7x-7-6x=-7

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now collect all your like terms on the left.

OpenStudy (anonymous):

\[(-6x-7x)+(6-7)=-7\]what do you get?

OpenStudy (anonymous):

-13x-1=-7

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now we need to get rid of the -1 on the left in order to isolate x. To do this, we add 1 to both sides. let me know what you get.

OpenStudy (anonymous):

-13x=-6

OpenStudy (anonymous):

yep. now solve for x. divide both sides by -13.

OpenStudy (anonymous):

wait we missed a term

OpenStudy (anonymous):

\[(-6x-7x-6x)+(6-7)=-7\]we forgot the 6x from earlier.

OpenStudy (anonymous):

Oh, I did that, haha I was wondering why I was confused!

OpenStudy (anonymous):

so we have -19x=-6

OpenStudy (anonymous):

now divide out the -19 on both sides. sorry for the confusion :P

OpenStudy (anonymous):

no worries (: 6/19

OpenStudy (anonymous):

yeah, looks good now.

OpenStudy (anonymous):

I'll triple check

OpenStudy (anonymous):

yeah...it's good.

OpenStudy (anonymous):

lol, thank you!!

OpenStudy (anonymous):

I also have another question

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Okay, for this equation: \[5\left[ 7x-4-3(x+1) \right]=-7x-2\] would I distribute the -3 first then distribute the 5?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[5(7x-4-3x-3)=-7x-2\]

OpenStudy (anonymous):

then distribute the 5. collect like terms. isolate x. solve for x.

OpenStudy (anonymous):

Okay !Thank you!

OpenStudy (anonymous):

I got 33/27 did I do something wrong?

OpenStudy (anonymous):

no that's what I get too

OpenStudy (anonymous):

you can simplify the fraction a bit. \[\frac{33}{27}=\frac{11}{9}\]

OpenStudy (anonymous):

Yep!! :D Thanks so much!

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

you're* typo...lol